Determine the molar solubility ( ? ) of Zn(CN)2 in a solution with a pH=5.89 . Ignore activities. The ?sp for Zn(CN)2 is 3.0×10−16 . The ?a for HCN is 6.2×10−10.
Answer #1

10 Kend Dale Foce HONT HO - CN + Hot - o ta tn pH=5. 89 x=(4+] = 10-st = 129 x 10-6 ko = [ca] [Hot] (HCN] .6.2x10-10 2 x² G-1

if you like my answer then plzzzz give positive ratings other wise ask you doubt in comment section I will solve it